What is the pH of a buffer that is 0.379 M HF and 0.374 M LiF? The Ka for HF is 3.5 x 10-4. If the ΔH = 109 kJ/mol and ΔS = 100 J/K mol, at what temperature does this reaction become spontaneous?
Answer #1

By using Henderson -- Hasselbalch Equation PH = Pka + log TAB Given – HF = 0.379 M. Lif = 0.374 M 7.5 X 0-4 Ka = 00374 0.379ANT 0 = 109 kJ/mol - TX 0-1 KJ/K TX 0-1 = 109 109 T= 1090 K, TE 87% Thank hobe this will ill help youtherefore the reaction becomes spontaneous when T = 1090K ( 817ºC)